How do I find the maximum value of Y=x^2-4 when working with quadratic functions
Answer by Alan3354(69443)

How do I find the maximum value of Y=x^2-4
-----------
That doesn't have a max. It has a minimum.
----
Find the vertex.
x of the vertex = -b/2a = 0/2
x = 0 --> y = -4
Minimum at (0,-4)